OpenStudy (anonymous):

Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? f(n) = 28(0.5)n f(n) = 56(0.5)n-1 f(n) = 56(0.5)n f(n) = 112(0.5)n-1

OpenStudy (anonymous):

the first one is 2

OpenStudy (dan815):

barry needs to get a life lol

OpenStudy (anonymous):

haha

OpenStudy (dan815):

can you see how the geese keep getting halved everyday

OpenStudy (anonymous):

its only one question so the answer is b?

OpenStudy (anonymous):

yes its B

OpenStudy (anonymous):

EXPLAIN

OpenStudy (anonymous):

ok wait a sec

OpenStudy (taramgrant0543664):

Doesn't have to be c?

OpenStudy (anonymous):

yeah im leaning toward c but im not sure yet

OpenStudy (taramgrant0543664):

I am assuming that n-1 should actually be in brackets like (n-1) then when f(1) then you get 0 where if you use c then it comes out as 56

OpenStudy (anonymous):

slade prolly got the answer from someone else thats the only thing i can think

OpenStudy (anonymous):

yes i should be in brackets so C is true

OpenStudy (anonymous):

no it didnt have brackets so i got confused and picked the next one that i thout

OpenStudy (ybarrap):

$$ f(1)=56\\ f(2)=28=\cfrac{f(1)}{2}\\\\ f(3)=14=\cfrac{f(2)}{2}=\cfrac{f(1)}{2^2}\\ f(4)=7=\cfrac{f(3)}{2}=\cfrac{f(1)}{2^3}\\ \cdots\\ f(n)=\cfrac{f(1)}{2^{(n-1)}}=\cfrac{56}{2^{(n-1)}}=56\left(\cfrac{1}{2}\right)^{n-1} $$ Does this make sense?